OpenStudy (anonymous):

write the slope of the equation of the line through the given points... through: (-2,3) and (1,5)

OpenStudy (anonymous):

Use the point-slope form. \[(y-y _{1})=m(x-x _{1)}\]

OpenStudy (anonymous):

the answer???/

OpenStudy (anonymous):

If you were to solve it, it would look like 5-3/-2-1= 2/-3

OpenStudy (anonymous):

The slope is 2/-3

OpenStudy (anonymous):

theres a second step to it isn't there

OpenStudy (anonymous):

I did the whole thing. If you want the whole answer, it's (5,3)=2/-3(1,-2)

OpenStudy (anonymous):

At least that's what my teacher wants from us

OpenStudy (anonymous):

my teacher wants the second part to it

OpenStudy (anonymous):

That's literally all there is. I don't know what else to say.

OpenStudy (anonymous):

You plug your number into the equation \[(y-y _{1})=m(x-x _{1})\], solve, which I did above, then put your numbers back into the equation with m being the slope

OpenStudy (anonymous):

oh really ^_^ so your saying this is the answer right;(5,3)=2/-3(1,-2)

OpenStudy (anonymous):

Best to my knowledge, yes

OpenStudy (anonymous):

thank you :)

OpenStudy (anonymous):

what about (1,2) and (4,3)

OpenStudy (anonymous):